# High-Precision Dichotomous Image Segmentation via Probing Diffusion Capacity

### v1

@@ Abstract
In the realm of high-resolution (HR), fine-grained image segmentation, the primary challenge is balancing broad contextual awareness with the precision required for detailed object delineation, capturing intricate details and the finest edges of objects. Diffusion models, trained on vast datasets comprising billions of image-text pairs, such as SD V2.1, have revolutionized text-to-image synthesis by delivering exceptional quality, fine detail resolution, and strong contextual awareness, making them an attractive solution for high-resolution image segmentation. To this end, we propose DiffDIS, a diffusion-driven segmentation model that taps into the potential of the pre-trained U-Net within diffusion models, specifically designed for high-resolution, fine-grained object segmentation. By leveraging the robust generalization capabilities and rich, versatile image representation prior of the SD models, coupled with a task-specific stable one-step denoising approach, we significantly reduce the inference time while preserving high-fidelity, detailed generation. Additionally, we introduce an auxiliary edge generation task to not only enhance the preservation of fine details of the object boundaries, but reconcile the probabilistic nature of diffusion with the deterministic demands of segmentation. With these refined strategies in place, DiffDIS serves as a rapid object mask generation model, specifically optimized for generating detailed binary maps at high resolutions, while demonstrating impressive accuracy and swift processing. Experiments on the DIS5K dataset demonstrate the superiority of DiffDIS, achieving state-of-the-art results through a streamlined inference process. Our code will be made publicly available.
@@ 1 Introduction
High-accuracy dichotomous image segmentation (DIS) (Qin et al., 2022) aims to accurately identify category-agnostic foreground objects within natural scenes, which is fundamental for a wide range of scene understanding applications, including AR/VR applications (Tian et al., 2022; Qin et al., 2021), image editing (Goferman et al., 2011), and 3D shape reconstruction (Liu et al., 2021). Different from existing segmentation tasks, DIS focuses on challenging high-resolution (HR) fine-grained object segmentation, regardless of their characteristics. These objects often encompass a greater volume of information and exhibit richer detail, thereby demanding more refined feature selection and more sophisticated algorithms for segmentation. Current CNN-(Pei et al., 2023; Qin et al., 2022) and Transformer-based (Xie et al., 2022; Kim et al., 2022; Yu et al., 2024) methods, despite their robust feature extraction capabilities, often face challenges in balance receptive field expansion with detail preservation in high-resolution images (Xie et al., 2022; Yu et al., 2024).
@@
However, there are several challenges when leveraging the diffusion model to DIS: On the one hand, for diffusion models: (1) The inherent high-dimensional, continuous feature space of DPM conflicts with the discrete nature of binary segmentation, potentially leading to a discrepancy in the predictive process. (2) Moreover, the diffusion models often suffer from lengthy inference time. Due to the recurrent nature of diffusion models, it usually takes more than 100 steps for DPM to generate satisfying results (Ke et al., 2024), which further exacerbates the already slow inference speeds of HR images due to their substantial data volume. (3) There is a fundamental conflict between the stochastic nature of diffusion and the deterministic outcomes required for image perception tasks. On the other hand, for the DIS task: As image resolution increases, so does the intricacy of the visual information and the corresponding computational demands, necessitating more refined feature selection.
@@
Against this backdrop, we propose DiffDIS leveraging the following strategies: First, we found that the Variational Autoencoder (VAE) (Kingma, 2013) has demonstrated the capability to nearly achieve perfect reconstruction of binary masks (See Tab. 1) . Therefore, following SD (Rombach et al., 2022), by mapping the masks through the VAE into the latent space, we can not only effectively leverage the strengths of diffusion models for denoising and refinement within it, but also significantly reduces the computational cost associated with processing high-accuracy HR image segmentation. Then, by employing a direct one-step denoising paradigm (See Eqn. 3) and integrating the pretrained parameters of SD-Turbo (Sauer et al., 2023; Parmar et al., 2024), which feature a smoother probability curve compared to the standard SD V2.1 parameters, we aim to streamline the network into an end-to-end model. This approach facilitates one-step denoising while leveraging its robust generalization capabilities and preserving the intricate details of high-resolution objects. Moreover, we introduce a joint predicting strategy for mask and edge, where the edge task assists in the training phase. This not only enhances the controllability of fine details but mitigates the misalignment between the generative nature of diffusion models and the determinism required in perception tasks. Finally, we utilize RGB latent representations as conditions for the diffusion model and introduce Scale-Wise Conditional Injection to establish multi-granular, long-range, and profound visual interactions.
In this paper, we successfully trained a one-step SD model that generates high-resolution, high-quality results in DIS, preserving fine details while accurately segmenting complex structures (See Fig. 1) . Generally, our main contributions can be summarized as follows:
@@
  • We significantly advance the field forward by outperforming almost all metrics on the DIS benchmark dataset, and thus establish a new SoTA in this space. Additionally, it boasts an inference speed that is orders of magnitude faster than traditional multi-step diffusion approaches, without compromising accuracy, demonstrating the superiority of DiffDIS.
@@ 2.1 Conventional Approach to Deal With DIS Works
Dichotomous Image Segmentation (DIS) is formulated as a category-agnostic task that focuses on accurately segmenting objects with varying structural complexities, independent of their specific characteristics. It includes high-resolution images of camouflaged (Fan et al., 2020; Le et al., 2019), salient (Pang et al., 2020; Zhao et al., 2020), and meticulous (Liew et al., 2021; Yang et al., 2020) objects in various backgrounds. What distinguishes DIS from conventional segmentation tasks is the demand for highly precise object delineation, including the finest internal details of objects. Upon introducing the DIS dataset, Qin et al. also presented IS-Net (Qin et al., 2022), a model specifically crafted to address the DIS challenge, utilizing the U2Net and intermediate supervision to mitigate overfitting risks. PF-DIS (Zhou et al., 2023) utilized a frequency prior generator and feature harmonization module to identify fine-grained object boundaries in DIS. UDUN (Pei et al., 2023) proposed a unite-divide-unite scheme to disentangle the trunk and structure segmentation for high-accuracy DIS. InSPyReNet (Kim et al., 2022) was constructed to generate HR outputs with a multi-resolution pyramid blending at the testing stage. Recent approaches have incorporated multi-granularity cues that harness both global and local information to enhance detail fidelity and object localization accuracy. BiRefNet (Zheng et al., 2024) employed a bilateral reference strategy, leveraging patches of the original images at their native scales as internal references and harnessing gradient priors as external references. Recently, MVANet (Yu et al., 2024) modeled the DIS task as a multi-view object perception problem, leveraging the complementary localization and refinement among different views to process HR, fine-detail images. Despite their commendable performance, existing methods have not effectively balanced the semantic dispersion of HR targets within a limited receptive field with the loss of high-precision details associated with a larger receptive field when tackling DIS. Compared to conventional approaches, our utilization of a diffusion architecture has excelled in achieving high-quality background removal, with fast processing times (0.3-0.4s), and offers a straightforward integration.
@@ 2.2 Diffusion Models for Dense Prediction and Efficient Inference in Diffusion
However, traditional multi-step generation models encounter a challenge: the recurrent structure of diffusion models often necessitates over 100 steps to produce satisfactory outputs. In response, various proposals have emerged to expedite the inference process. Distillation-based methods have demonstrated remarkable speedups by optimizing the original diffusion model’s weights with enhanced schedulers or architectures.  (Luo et al., 2023) achieves a few-step inference in conditional image generation through self-consistency enforcement. SD-Turbo (Sauer et al., 2023) employs Adversarial Diffusion Distillation (ADD), utilizing a pre-trained SD model to denoise images and calculate adversarial and distillation losses, facilitating rapid, high-quality generation. Recently, GenPercept (Xu et al., 2024) presents a novel perspective on the diffusion process as an interpolation between RGB images and perceptual targets, effectively harnessing the pre-trained U-Net for various downstream image understanding tasks. Given that SD-Turbo refines both high-frequency and low-frequency information through distillation (Sauer et al., 2023), by harnessing its efficient prior for image generation, we can sustain competitive performance in a single-step denoising scenario while achieving high-fidelity, fine-grained mask generation.
@@ 4.2 Edge-Assisted Training Strategy
In high-resolution tasks, the complexity and detail richness make it tough to capture fine features. Furthermore, for diffusion-based architectures, accurately depicting detailed features during few-step generation poses a challenge, especially as details may be overlooked during denoising in the latent space, which is often reduced to $\frac{1}{8}$ the size of the input. Some studies (Pei et al., 2023; Chen et al., 2023; Cheng et al., 2020) have indicated that incorporating additional edge constraints improves the boundary segmentation performance of masks. Given these considerations, we propose an integrated and streamlined network architecture that concurrently predicts noise for both the mask and edge, leveraging auxiliary edge information to constrain the network. These dual prediction streams operate within the same network structure and share parameters, with batch discriminative embedding applied to distinguish between them effectively. Within the U-Net architecture, specifically in the mid-block where semantic information is most concentrated, we incorporate a task-specific enhancement by upgrading the original attention module to a Detail-Balancing Interactive Attention. This mechanism aligns the attention regions of both the mask and edge streams, facilitating more efficient interaction and complementarity between the two.
@@
Our method differs from the approach of (Fu et al., 2024) in how we handle feature interaction. In their work, the keys ($\mathbf{K}$) and values ($\mathbf{V}$) from different domains are concatenated into a unified representation, and cross-attention is performed by applying the queries ($\mathbf{Q}$) from each domain to the concatenated $\mathbf{K}$ and $\mathbf{V}$. Our method surpasses theirs by enabling direct, targeted attention operations that facilitate a more nuanced interaction between mask and edge features, leading to better discrimination and alignment without the dilution of information (See in Tab. 5) .
@@ 4.3 Scale-Wise Conditional Injection
In traditional diffusion models for dense prediction tasks, conditioning is typically done at the input stage by concatenating along the channel dimension with the variable to be denoised (Ke et al., 2024; Fu et al., 2024). This approach can lead to information loss in later stages. To establish long-range and profound conditional guidance, we introduce Scale-Wise Conditional Injection to create multi-granular, long-range, and deep visual interactions. Specifically, we incorporate multi-scale conditions into the corresponding layers of the U-Net encoder (See Fig. 3) . Specifically, we employ three injector heads, each composed of a simple convolutional layer and two zero conloution layers (Zhang et al., 2023) (See Fig. 3). Each injector head receives conditional latent code that is resized to corresponding scales as input. The resulting outputs are then integrated at the junction of the last three layers in the U-Net encoder, facilitating the generation of more authentic structural details. In this way, we introduce multi-granularity information for semantic and structural perception, without causing excessive interference with RGB texture.
@@ 5.1 Experimental Settings
Implementation Details Experiments were implemented in PyTorch and conducted on a single NVIDIA H800 GPU. During training, the original images were resized to $1024\times 1024$ for training. We use SD V2.1 (Rombach et al., 2022) as the backbone, and initialize the model with the parameters from SD-Turbo (Sauer et al., 2023). To mitigate the risk of overfitting in diffusion models when trained on a relatively small dataset, we apply several data augmentation techniques, including random horizontal flipping, cropping, rotation, and CutMix (Yun et al., 2019). For optimization, we use the Adam optimizer, setting the initial learning rate to $3\times 10^{-5}$. The batch size is configured as $2$. The maximum number of training epochs was set to $90$. During evaluation, we binarize the predicted maps for accuracy calculation.
@@ 5.2 Comparison
DIS-VD $F_{\beta}^{max}$uparrow 0.738 0.763 0.749 0.798 0.791 0.823 0.823 0.889 0.897 0.913 0.844 0.918
$F^{\omega}_{\beta}$uparrow 0.615 0.691 0.653 0.733 0.717 0.763 0.763 0.834 0.863 0.856 0.824 0.888
$E^{m}_{\phi}$uparrow 0.807 0.852 0.829 0.879 0.856 0.891 0.892 0.914 0.937 0.938 0.924 0.948
@@
$\mathcal{M}$downarrow 0.100 0.080 0.088 0.067 0.074 0.062 0.059 0.042 0.036 0.036 0.044 0.029
DIS-TE1 $F_{\beta}^{max}$uparrow 0.683 0.717 0.673 0.754 0.740 0.784 0.784 0.845 0.866 0.893 0.807 0.903
@@
Overall $F_{\beta}^{max}$uparrow 0.744 0.778 0.755 0.809 0.799 0.831 0.831 0.891 0.900 0.916 0.840 0.921
$F^{\omega}_{\beta}$uparrow 0.618 0.707 0.661 0.746 0.726 0.770 0.772 0.838 0.865 0.855 0.817 0.892
$E^{m}_{\phi}$uparrow 0.811 0.864 0.832 0.885 0.858 0.895 0.892 0.917 0.938 0.938 0.920 0.947
$S_{m}$uparrow 0.789 0.812 0.791 0.830 0.819 0.847 0.844 0.900 0.906 0.905 0.860 0.905
@@ 5.3 Ablation
0.895 0.916 0.880 0.046
$\checkmark$ 0.902 0.921 0.888 0.036
$\checkmark$ $\checkmark$ 0.917 0.943 0.895 0.032
$\checkmark$ $\checkmark$ $\checkmark$ 0.918 0.945 0.899 0.030
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.918 0.948 0.904 0.029
@@
Train from scratch 1 1 0.747 0.758 0.687 0.103 0.33
SD-Turbo 1 1 0.895 0.916 0.880 0.046 0.33
1* 2 0.889 0.913 0.878 0.047 0.52
2 2 0.894 0.923 0.881 0.038 0.52
SDV2.1 1 1 0.890 0.910 0.869 0.048 0.33
1* 10 0.823 0.878 0.854 0.052 1.40
@@
The last two rows of Table 4 demonstrate the comparative efficacy between one-step denoising paradigms and traditional multi-step approaches (Ke et al., 2024; Rombach et al., 2022) in the context of high-precision, HR object segmentation. The final row, which employs a random timestep between 0 and 1000 to introduce noise into the mask latent during training, followed by a ten-step denoising process during testing, similar to (Ke et al., 2024), yields significantly lower results than one-step paradigms. This diminished performance is attributed to the increased prediction variance stemming from the model’s exposure to inconsistent data distributions during training.
@@
$\checkmark$ $\checkmark$ 0.917 0.943 0.895 0.032
$\checkmark$ $\checkmark$ 0.912 0.935 0.894 0.035

### v2

@@ Abstract
In the realm of high-resolution (HR), fine-grained image segmentation, the primary challenge is balancing broad contextual awareness with the precision required for detailed object delineation, capturing intricate details and the finest edges of objects. Diffusion models, trained on vast datasets comprising billions of image-text pairs, such as SD V2.1, have revolutionized text-to-image synthesis by delivering exceptional quality, fine detail resolution, and strong contextual awareness, making them an attractive solution for high-resolution image segmentation. To this end, we propose DiffDIS, a diffusion-driven segmentation model that taps into the potential of the pre-trained U-Net within diffusion models, specifically designed for high-resolution, fine-grained object segmentation. By leveraging the robust generalization capabilities and rich, versatile image representation prior of the SD models, coupled with a task-specific stable one-step denoising approach, we significantly reduce the inference time while preserving high-fidelity, detailed generation. Additionally, we introduce an auxiliary edge generation task to not only enhance the preservation of fine details of the object boundaries, but reconcile the probabilistic nature of diffusion with the deterministic demands of segmentation. With these refined strategies in place, DiffDIS serves as a rapid object mask generation model, specifically optimized for generating detailed binary maps at high resolutions, while demonstrating impressive accuracy and swift processing. Experiments on the DIS5K dataset demonstrate the superiority of DiffDIS, achieving state-of-the-art results through a streamlined inference process. The source code will be publicly available at DiffDIS.
@@ 1 Introduction
High-accuracy dichotomous image segmentation (DIS) (Qin et al., 2022) aims to accurately identify category-agnostic foreground objects in natural scenes, which is fundamental for a wide range of scene understanding applications, including AR/VR applications (Tian et al., 2022; Qin et al., 2021), image editing (Goferman et al., 2011), and 3D shape reconstruction (Liu et al., 2021). Different from existing segmentation tasks, DIS focuses on challenging high-resolution (HR) fine-grained object segmentation with various characteristics. These objects often encompass a greater volume of information and exhibit richer detail, thereby demanding more refined feature selection and more sophisticated algorithms for segmentation.
Current CNN-(Pei et al., 2023; Qin et al., 2022) and Transformer-based (Xie et al., 2022; Kim et al., 2022; Yu et al., 2024) methods, despite their robust feature extraction capabilities, often face challenges in balance receptive field expansion with detail preservation in high-resolution images (Xie et al., 2022; Yu et al., 2024).
@@
However, there are several challenges when leveraging the diffusion model to DIS: (1) The inherent high-dimensional, continuous feature space of DPM conflicts with the discrete nature of binary segmentation, potentially leading to a discrepancy in the predictive process. (2) Moreover, the diffusion models often suffer from lengthy inference time. Due to the recurrent nature of diffusion models, it usually takes more than 100 steps for DPM to generate satisfying results (Ke et al., 2024), which further exacerbates the already slow inference speeds of HR images due to their substantial data volume. (3) There is a fundamental conflict between the stochastic nature of diffusion and the deterministic outcomes required for image perception tasks.
@@
Against this backdrop, we propose DiffDIS, addressing both aforementioned challenges and task-specific complexities of DIS, focusing on enhanced processing speed, stronger detail perception, and higher determinism, achieved through the following strategies: First, we found that the Variational Autoencoder (VAE) (Kingma, 2013) has demonstrated the capability to nearly achieve perfect reconstruction of binary masks (See Tab. 1) . Therefore, following SD (Rombach et al., 2022), by mapping the masks through the VAE into the latent space, we can not only effectively leverage the strengths of diffusion models for denoising and refinement within it, but also significantly reduces the computational cost associated with processing high-accuracy HR image segmentation. Then, by employing a direct one-step denoising paradigm (See Eqn. 3) and integrating the pretrained parameters of SD-Turbo (Sauer et al., 2023; Parmar et al., 2024), which feature a smoother probability curve compared to the standard SD V2.1 parameters, we aim to streamline the network into an end-to-end model. This approach facilitates one-step denoising while leveraging its robust generalization capabilities and preserving the intricate details of high-resolution objects. Moreover, we introduce a joint predicting strategy for mask and edge , which enhances mask generation through finer constraints while improving controllability of diffusion models for perception tasks. Finally, we condition the diffusion model on RGB latent representations and propose Scale-Wise Conditional Injection to enable multi-granular, long-range feature interactions, ensuring more refined feature preservation and selection.
Generally, our main contributions can be summarized as follows:
@@
  • We advance the field forward by outperforming almost all metrics on the DIS benchmark dataset, and thus establish a new SoTA in this space. Additionally, it boasts an inference speed that is orders of magnitude faster than traditional multi-step diffusion approaches without compromising accuracy.
@@ 2.1 Conventional Approach to Deal With DIS Works
Dichotomous Image Segmentation (DIS) is formulated as a category-agnostic task that focuses on accurately segmenting objects with varying structural complexities, independent of their specific characteristics. It includes high-resolution images containing salient (Pang et al., 2020; Zhao et al., 2020; 2024c), camouflaged (Pang et al., 2024b; 2022; Fan et al., 2020; Pang et al., 2024a), and meticulous (Liew et al., 2021; Yang et al., 2020) instances in various backgrounds, integrating several context-dependent (Zhao et al., 2024b; a) segmentation tasks into a unified benchmark. When dealing with DIS, it is essential to consider the demand for highly precise object delineation, capturing the finest internal details of objects. Upon introducing the DIS dataset, Qin et al. also presented IS-Net (Qin et al., 2022), a model specifically crafted to address the DIS challenge, utilizing the U2Net and intermediate supervision to mitigate overfitting risks. PF-DIS (Zhou et al., 2023) utilized a frequency prior generator and feature harmonization module to identify fine-grained object boundaries in DIS. UDUN (Pei et al., 2023) proposed a unite-divide-unite scheme to disentangle the trunk and structure segmentation for high-accuracy DIS. InSPyReNet (Kim et al., 2022) was constructed to generate HR outputs with a multi-resolution pyramid blending at the testing stage. Recent approaches have incorporated multi-granularity cues that harness both global and local information to enhance detail fidelity and object localization accuracy. BiRefNet (Zheng et al., 2024) employed a bilateral reference strategy, leveraging patches of the original images at their native scales as internal references and harnessing gradient priors as external references. Recently, MVANet (Yu et al., 2024) modeled the DIS task as a multi-view object perception problem, leveraging the complementary localization and refinement among different views to process HR, fine-detail images. Despite their commendable performance, existing methods have not effectively balanced the semantic dispersion of HR targets within a limited receptive field with the loss of high-precision details associated with a larger receptive field when tackling DIS. Compared to conventional approaches, our utilization of a diffusion architecture has excelled in achieving high-quality background removal, with fast processing times, and offers a straightforward integration.
@@ 2.2 Diffusion Models for Dense Prediction and Efficient Inference in Diffusion
However, traditional multi-step generation models encounter a challenge: the recurrent structure of diffusion models often necessitates over 100 steps to produce satisfactory outputs. In response, various proposals have emerged to expedite the inference process. Distillation-based methods have demonstrated remarkable speedups by optimizing the original diffusion model’s weights with enhanced schedulers or architectures.  (Luo et al., 2023) achieves a few-step inference in conditional image generation through self-consistency enforcement. SD-Turbo (Sauer et al., 2023) employs Adversarial Diffusion Distillation (ADD), utilizing a pre-trained SD model to denoise images and calculate adversarial and distillation losses, facilitating rapid, high-quality generation. Recently, GenPercept (Xu et al., 2024) presents a novel perspective on the diffusion process as an interpolation between RGB images and perceptual targets, effectively harnessing the pre-trained U-Net for various downstream image understanding tasks. Given that SD-Turbo refines both high-frequency and low-frequency information through distillation (Sauer et al., 2023), by harnessing its efficient prior for image generation, we can sustain competitive performance in a single-step denoising scenario while achieving high-fidelity, fine-grained mask generation (See Fig. 1).
@@ 4.2 Edge-Assisted Training Strategy
In high-resolution tasks, the complexity and detail richness make it tough to capture fine features. Furthermore, for diffusion-based architectures, accurately depicting detailed features during few-step generation poses a challenge, especially as details may be overlooked during denoising in the latent space, which is often reduced to $\frac{1}{8}$ the size of the input. Some studies (Pei et al., 2023; Chen et al., 2023; Zhao et al., 2022) have indicated that incorporating additional edge constraints improves the boundary segmentation performance of masks. Given these considerations, we propose an integrated and streamlined network architecture that concurrently predicts noise for both the mask and edge, leveraging auxiliary edge information to constrain the network. These dual prediction streams operate within the same network structure and share parameters, with batch discriminative embedding applied to distinguish between them effectively. Within the U-Net architecture, specifically in the mid-block where semantic information is most concentrated, we incorporate a task-specific enhancement by upgrading the original attention module to a Detail-Balancing Interactive Attention. This mechanism aligns the attention regions of both the mask and edge streams, facilitating more efficient interaction and complementarity between the two.
@@
In their work, keys ($\mathbf{K}$) and values ($\mathbf{V}$) are concatenated into a unified representation, and cross-attention is computed using domain-specific queries ($\mathbf{Q}$) on this combined feature. Our method performs targeted attention between mask and edge features, enhancing discrimination and alignment while preserving information integrity (Tab. 5).
@@ 4.3 Scale-Wise Conditional Injection
In traditional diffusion models for dense prediction tasks, conditioning is typically done at the input stage by concatenating along the channel dimension with the variable to be denoised (Ke et al., 2024; Fu et al., 2024). This approach can lead to information loss in later stages. To establish long-range and profound conditional guidance, we introduce Scale-Wise Conditional Injection (SWCI), an approach that employs mixed-scale architectures to effectively capture diverse visual patterns, thereby enhancing multi-granular perception and facilitating deep visual interactions. Specifically, we incorporate multi-scale conditions into the corresponding layers of the U-Net encoder (See Fig. 3) and employ three injector heads, each composed of a simple convolution layer and two zero convolution layers (Zhang et al., 2023) (See Fig. 3). Each injector head receives conditional latent code that is resized to corresponding scales as input. The resulting outputs are then integrated at the junction of the last three layers in the U-Net encoder, facilitating the generation of more authentic structural details. In this way, we introduce multi-granularity information for semantic and structural perception, without causing excessive interference with RGB texture.
@@ 5.1 Experimental Settings
Implementation Details Experiments were implemented in PyTorch and conducted on a single NVIDIA H800 GPU. During training, the original images were resized to $1024\times 1024$ for training. We use SD V2.1 (Rombach et al., 2022) as our backbone, and initialize the model with the parameters from SD-Turbo (Sauer et al., 2023). To mitigate the risk of overfitting in diffusion models when trained on a relatively small dataset, we apply several data augmentation techniques, including random horizontal flipping, cropping, rotation, and CutMix (Yun et al., 2019). For optimization, we use the Adam optimizer, setting the initial learning rate to $3\times 10^{-5}$. The batch size is configured as $4$. The maximum number of training epochs was set to $90$. During evaluation, we binarize the predicted maps to filter out minor noise for accuracy calculation.
@@ 5.2 Comparison
DIS-VD $F_{\beta}^{max}$uparrow 0.738 0.763 0.749 0.798 0.791 0.823 0.823 0.889 0.897 0.904 0.844 0.908
$F^{\omega}_{\beta}$uparrow 0.615 0.691 0.653 0.733 0.717 0.763 0.763 0.834 0.863 0.863 0.824 0.888
$E^{m}_{\phi}$uparrow 0.807 0.852 0.829 0.879 0.856 0.891 0.892 0.914 0.937 0.941 0.924 0.948
@@
$\mathcal{M}$downarrow 0.100 0.080 0.088 0.067 0.074 0.062 0.059 0.042 0.036 0.034 0.044 0.029
DIS-TE1 $F_{\beta}^{max}$uparrow 0.683 0.717 0.673 0.754 0.740 0.784 0.784 0.845 0.866 0.873 0.807 0.883
@@
Overall DIS-TE(1-4) $F_{\beta}^{max}$uparrow 0.744 0.778 0.755 0.809 0.799 0.831 0.831 0.891 0.900 0.908 0.840 0.911
$F^{\omega}_{\beta}$uparrow 0.618 0.707 0.661 0.746 0.726 0.770 0.772 0.838 0.865 0.861 0.817 0.892
$E^{m}_{\phi}$uparrow 0.811 0.864 0.832 0.885 0.858 0.895 0.892 0.917 0.938 0.938 0.920 0.950
$S_{m}$uparrow 0.789 0.812 0.791 0.830 0.819 0.847 0.844 0.900 0.906 0.904 0.860 0.905
@@ 5.3 Ablation
0.882 0.930 0.880 0.039
$\checkmark$ 0.891 0.934 0.888 0.036
$\checkmark$ $\checkmark$ 0.896 0.943 0.895 0.032
$\checkmark$ $\checkmark$ $\checkmark$ 0.897 0.945 0.899 0.030
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.908 0.948 0.904 0.029
@@
Train from scratch 1 1 0.682 0.758 0.687 0.103 0.33
SD-Turbo 1 1 0.883 0.930 0.880 0.039 0.33
1* 2 0.876 0.918 0.878 0.041 0.52
2 2 0.879 0.933 0.881 0.038 0.52
SDV2.1 1 1 0.853 0.910 0.862 0.048 0.33
1* 10 0.837 0.888 0.831 0.052 1.40
@@
The last two rows of Table 4 demonstrate the comparative efficacy between one-step denoising paradigms and traditional multi-step approaches (Ke et al., 2024; Rombach et al., 2022) in the context of high-precision, HR object segmentation. The final row, which employs a random timestep between 0 and 1000 to introduce noise into the mask latent during training, followed by a ten-step denoising process during testing, similar to (Ke et al., 2024), yields significantly lower results than one-step paradigms. This diminished performance is attributed to the randomness of multi-step denoising, where stochastic variations are progressively amplified at each step. This leads to increased unpredictability and generative characteristics, conflicting with the deterministic nature of binary segmentation tasks that require unambiguous outputs.
@@
$\checkmark$ $\checkmark$ 0.896 0.943 0.895 0.032
$\checkmark$ $\checkmark$ 0.891 0.940 0.894 0.035